Instrument Log

 

USDA UVB   C1  Lamont, OK
 

Date

Technician

Comments

4/19/16

KT

Instrument mentor could not communicate with the visible instrument. At 1355 GMT observed that the shadow band was not rotating. Power cycled the mother board at 1400 GMT. Instrument mentor was able to connect to the instrument. RTS at 1410 GMT.

10/14/14

CAM

Instrument mentor notified Site Ops he was unable to communicate with the VIS and UV instruments.  Checked comms locally and was able to communicate with both data loggers by connecting directly to the board and through the serial cable back at the MOXAs.  Further T/S revealed  both MOXAs had failed.  IT department replaced both MOXAs at 1715 GMT.  Mentor confirmed remote comms were restored and UV data looked good, but all VIS channels are reading the same.  Measured the outputs of the gain resistors for channels 1-7 on the logger board and confirmed all were the same, 1.7V.  Power cycled the board and checked voltages again.  This time voltages read different as expected and varied with covering/uncovering of sensor.  RTS status O.K. at 1945 GMT.

8/27/14

CAM

Instrument mentor sent 2 Black Box SP 361A-R2 RS232 surge protectors and asked that they be installed.  Installed surge protectors on both data logger serial ports at 1230 GMT.

8/19/14

CAM

Instrument mentor notified site ops that he was unable to communicate with the logger boards.  Troubleshooting found that the RS-232 chips on both the VIS and UV logger boards had failed again.  Replaced the MAX231 chips and local communications were restored at 1930 GMT.  Notified mentor that instruments were ready to be restarted.

7/1/14

CAM

Bill Durham notified Site Ops he was unable to communicate with the VIS datalogger.  T/S found the MAX231 RS232 chip had failed on the logger board.  R/R MAX231 chip at 1250 GMT and verified comms locally. Left voice mail with mentor that instrument was ready to be started.

10/8/13

JEM

An attempted communication upgrade caused collections failures and resulted in data loss from approximately 1700 GMT on 10/8 to an unknown time on 10/11 GMT.

9/17/13

CAM

George Janson notified Site Ops he was unable to communicate with the VIS datalogger.  T/S found the MAX231 RS232 chip had failed on the logger board.  R/R MAX231 chip and verified comms locally. Contacted mentor who was then able to communicate remotely.  RTS at 1550 GMT.  Note there were thunderstorms in the area the previous day and night.

6/28/13

JEM

Instrument mentor notified site ops that he was unable to communicate with the logger boards again.  Troubleshooting found that the RS-232 chips on the logger boards had failed again.  Replaced the chips and local communications were restored.  Notified mentor and he logged in, but was still unable to communicate with the logger boards.  Moved the VIS instrument to port 5 and the UV instrument to port 7.  Mentor retested communications and everything was OK.  RTS @2030 GMT.

6/26/13

CAM

Instrument mentor (George Janson) notified Site Ops that he was unable to communicate with either the VIS or UV instrument.  T/S of VIS instrument found the 1.5A board fuse was blown.  R/R fuse but still could not communicate with logger.  Further T/S found the MAX201 RS232 communications chip had failed.  R/R chip and could then communicate with logger when connected directly to it but could not communicate through the Ethernet-serial adapters.  The Ethernet-serial adapters at both ends had failed.  R/R adapters and could then communicate with logger through the Ethernet line.  T/S of UV instrument found its MAX201 RS232 chip had also failed.  R/R chip and could then communicate with logger directly and through Ethernet-serial connection.  Contacted mentor who tried to communicate remotely but could not get through the Rocketport.  Further T/S found both ports 2 and 4 had failed on the Rocketport.   Moved the VIS instrument from port 2 to port 6 and the UV instrument from port 4 to port 8.  Mentor was then able to access instruments remotely.  Both instruments restarted and RTS at 2150 GMT.  Note a severe thunderstorm passed over at time of outage yesterday and other nearby network equipment also suffered damage.

4/29/13

CAM

The RH sensor has been reading well over 100% during recent rain events.  The previous ¼ turn adjustment to the dry potentiometer had no effect.  Mentor requests an additional full turn CCW.  Adjusted dry potentiometer 1 full turn CCW at 1620 GMT.

4/25/13

CAM

George Janson notified Site Ops that the USDA RH sensor was reading 109% during recent rain.  Requested Site Ops adjust the dry potentiometer ¼ turn CCW.  Made adjustment as requested at 2010 GMT.

4/23/13

CAM

George Janson from CSU notified Site Ops that channel 1 on the visible instrument was showing failure beginning in early April.   Powered datalogger off at 1530 GMT after data was remotely collected by CSU.  Removed head cable and observed light corrosion on Conxall cable end.  Cleaned cable pins and head sockets with contact cleaner and a pick.  Checked pins on Amphenol connector at board and all were secure.  Cleaned connector with contact cleaner.  Restored power at 1540 GMT.  Measured voltage before and after gain resistors for all 7 channels.  All voltages were reasonable.  George Janson restarted instrument and reported channel 1 problem was corrected.

3/7/13

CAM

Bill Durham from CSU requested the +5V power to the VIS air temperature system be cycled.  Disconnected +5V connector for 5 seconds.  Reconnected and RTS at 1820 GMT.

3/7/13

CAW

Received the repaired VIS head back from repair.  The diffuser had been damaged by birds and was replaced.   Need to install head.  Installed *VIS head at 1400 GMT.  Notified mentor.  *New - 220

2/28/13

CAM

George Janson notified Site Ops he was unable to communicate with the VIS datalogger.  T/S found the CAT5 cable had pulled out of the serial adapter at the Rocket Port.  Reconnected cable at 1830 GMT and verified communications were restored.

6/27/12

CAM

Bill Durham notified Site Ops he was unable to communicate with the UV instrument.  Check of instrument found shadow band was not rotating.  Connected laptop to data logger and was unable to communicate locally as well.  Power cycled data logger at 1410 GMT and successfully reestablished comms.  Notified Bill Durham who was then able to restart instrument.

5/21/12

PAD/KT

Diffuser on USDA VIS head has been punctured.  Mentor requests *VIS head removed and sent for repairs and calibration.  Shut off power and removed head at 1605 GMT.  Powered controller board up after head removal at mentor’s request.  RTS at 1620 GMT.   *Old - 220

5/1/12

CAM

Bill Durham from CSU notified Site Ops that communications with the VIS instrument were lost overnight.  There was severe weather in the area around the time of outage.  Check of instrument found shadow band was not rotating.  Power cycled logger board but still could not establish comms.  Replaced MAX201CPD RS232 chip and could then communicate with logger locally.  Notified mentor who was then able to remotely initialize and start the instrument.  RTS status O.K. at 2004 GMT.

4/30/12

CAM

Bill Durham from CSU notified Site Ops that communications with the UV instrument were lost sometime Friday night.  Check of instrument found shadow band was not rotating.  Power cycled logger board at 1630 GMT and notified mentor.  Mentor returned call at 1650 GMT and advised instrument was operational.  The heater is not working, but that was a pre-existing problem that mentor is aware of.  Confirmed band operation and proper shading at 1700 GMT.

4/10/12

PAD

Notified by mentor via email shadowband on UV instrument needs aligned.  At 1831 GMT, observed shadowband was not coming quite far enough to completely shade diffuser.  Adjusted shadowband and verified diffuser is now completely shading during measurement cycle.  RTS at 1846 GMT.

2/23/12

CAM

George Janson from CSU notified Site Ops that the head temperature on the UV instrument was not controlling.  He requested the head cable connections be cleaned.  Powered UV instrument off at 1945 GMT and disconnected head cable.  Observed a little corrosion on the cable and head connections, but nothing major.  Cleaned connections and restored power at 1952 GMT.  Notified mentor.

2/23/12

CAM

George Janson from CSU notified Site Ops that he was unable to communicate with either one of the USDA instruments.  He requested both boards be power cycled.  Observed both shadow bands rotating as they should be.  Power cycled both the UV and VIS boards at 1820 GMT.  Checked Rocket Port and found the instrument data cables had been inadvertently disconnected during unrelated maintenance the previous day.  Connected data cables to the correct ports and RTS at 1825 GMT.  Mentor was then able to communicate with both instruments.

7/5/11

CAM

Bill Durham from Colorado State notified Site Ops that he was unable to communicate with UV and VIS instruments.  T/S found neither shadow band was rotating, but all logger board test point voltages were good.  Connected to each logger directly and received no response from either.   Replaced the MAX231 comms chip on the VIS board and established comms at 1400 GMT.  The UV board originally had a MAX201 chip, but no spares were available.  Replaced the UV logger’s MAX201 with a MAX231 chip, as they are interchangeable.  Established comms directly with UV logger at 1430 GMT.  Notified Bill Durham who was then able to communicate remotely with the VIS instrument, but not the UV instrument.  Further T/S found port 3 on the Rocketport had failed.  Moved UV instrument to port 4 and notified Bill of change.  RTS at 1600 GMT.

6/23/11

CAM

Neither the VIS or UV instrument has been running since a lightning strike occurred on 6/12.  T/S found that both logger boards had failed.  George Janson from Colorado State sent replacement boards.  Installed new *VIS logger board at 1225 GMT and verified communication.  Installed new **UV logger board at 1236 and verified communication.  Notified Colorado State personnel that instruments were ready to be started and operational status confirmed.  *Old – 93931  *New – 93924  **Old – BD26  **New – CFD7

4/2/07

CAM

The shadow band on the Vis instrument is rotating continuously.  Mentor requested the motor IC chip be replaced.  Powered logger board off at 1900 GMT.  R/R motor chip.  Restored power at 1902 GMT.  Notified mentor, who restarted the instrument.  New chip did not correct the problem.  Mentor will send a replacement motor and cable.

1/16/07

MK

UV head needs to be replaced due to open heater. R/R *UV head at 2130 GMT, restarted at 2132.   *Old – 283   *New - 387

8/1/06

MK

Spare *logger board for the VIS-UVB system needs to be installed. R/R VIS board at 1330 GMT. Power on AC power, modem and VIS board. *Old - 93931   *New - 93932

7/28/06

MK

The UVB instruments are not communicating at this time. Due to the overnight power outages the UVB modems are not transmitting data. Replaced 2 amp fuse on the VIS board, reset modems and logger boards. All communications were restored with both instruments, however a problem has developed with the VIS board. After troubleshooting problem with George Jansen he has determined that the logger board could have been damaged from the storm and a replacement board is needed. He is sending a new logger board and it could be installed as soon as 8/1/06.

4/10/06

MK

The communications equipment needs to be moved and reorganized into a larger enclosure. The UPS and power supplies need to be moved and reorganized into a different enclosure. Disconnected power and communications to all instruments at E13 comms enclosure at 1525 GMT. Moved UPS and power supplies, organized all communications equipment and powered back on at 1945 GMT.

2/22/06

MK

Communications with UV-B Monitoring station are down at this time. The AA led (auto answer) on the modem is currently off. Power cycled modem at 1636 GMT. All modem LED's returned to normal.

3/24/05

MR

The new data logger has arrived and needs to be installed.  Installed the new data logger as per mentors instructions.  The correct values are now being reported.  RTS at 2115 GMT.

3/23/05

MR

The instrument mentor has contacted me and notified me that the instrument is still not reporting the correct values and wants further T/S performed.  Performed further T/S with the mentor on the phone.  It was determined that the data logger board needs to be replaced.  The mentor will send a new board out to us and wants it installed ASAP when it arrives.

3/23/05

CAM

Instrument mentor (Bill Durham) called and requested the visible instrument's logger board be powered off for 5 minutes and then back on.   The instrument has been down since the recent thunderstorms.  Powered the logger board off at 1440 GMT.  Powered back on at 1445 GMT.  Mentor advised he will call the instrument within the next 15 minutes to try and restart it.

3/22/05

MR

The instrument mentor wants the instrument power cycled due to no communications.  Power cycled the instrument.  RTS at 1445.

2/16/05

MR

The instrument mentor has sent a replacement head for the Visible sensor.  R/R the visible sensor at 1530 GMT.  Called the mentor at 1540 GMT and had him restart the instrument.  Instrument restarted at 1543 GMT status OK.  Returned the old head to the instrument mentor.

1/25/05

CAM

Instrument mentor advised that the visible instrument’s temperature sensor stopped reporting and requested that the sensor’s 5V power supply be cycled.  Disconnected the 5V power connector at 1530 GMT.  Waited a few seconds and reconnected.  Mentor then remotely logged in and verified the temperature sensor was working.  RTS status O.K.

1/20/05

MR

The instrument mentor is having trouble connecting to the UV instrument.  T/S with the mentor on the phone and found that the telephone modem switch is not switching over to the UV instrument as instructed.  The mentor will send another switch to us overnight.  RTS at 1645 GMT.

7/2/04

CAM

Instrument mentor requested that both the USDA visible and UV instruments be power cycled due to loss of communications.  Observed that the UV instrument was running, but the visible was not.  Powered both instruments off at 1413 GMT.  Waited 2 minutes and powered back on at 1415 GMT.  Called and left voice mail informing mentor of actions.  Mentor returned call at 1435 GMT and advised that both instruments were successfully restarted.

6/3/04

MR

The instrument mentor notified us that the USDA visible and infrared instrument needs to be reset.  Suspect they shut down after a major lightening strike near the CF on 6/2/04.  Called the mentor and he had me power cycle both data loggers.  He then was able to establish communications with the instruments and restart them.  RTS at 1415 GMT.

2/10/04

CAM

While checking the visible MFR alignment as requested by mentor, the UV MFR shadow band was observed out of position.  The band was all the way to the west and striking the bottom of the UV MFR housing.  The band was always trying to rotate in the same direction instead of switching directions to return to the home position.  Notified the instrument mentor who advised the phone switch had connected him to the wrong instrument while trying to align the visible MFR remotely.  Power cycled the UV MFR at 1845 GMT.  Mentor restarted the UV MFR and the shadow band was then properly aligned at 1850 GMT.  RTS status O.K. at 1852 GMT.

2/10/04

CAM

Instrument mentor requested that the visible MFR alignment be checked.  Found the alignment was slightly off, not completing shading the sensor.  Adjusted the band so that it was centered over the sensor.  RTS at 1750 GMT.

1/28/04

MR

The instrument mentor wants us to check the alignment of the UV shadow band.  Checked the shadow band alignment and found it was slightly off.  Adjusted the alignment.  Notified the mentor.  RTS at 1900 GMT.

1/8/04

MR

The infrared sensor needs to be releveled.  Leveled the sensor.  RTS at 2115 GMT.

8/5/03

MR

The instrument mentor wants the visible instrument to be reset.    Power cycled the data logger at 1330 GMT.  RTS at 1345 GMT.

5/21/03

MR

The instrument mentor has sent a new data logger board to replace the current one for the visible MFR. R/R the *data logger board.  Further T/S found the Temperature probe was also bad.  The mentor will send the parts to repair the temperature probe.  RTS at 1745 GMT.  *Old - B30C   *New - C148. 

4/11/03

MR

The instrument mentor has requested that the *MFR sensor head be changed out for calibration.  Called the instrument mentor at 1720 to allow him to collect the data.  Removed the old sensor at 1724 GMT.  Installed the new sensor and turned the power back on.  The instrument mentor then was able to connect to the instrument and restart it at 1733 GMT.  RTS at 1735 GMT.  *Old - 303   *New - 288

2/4/03

JEM

The entire Central Facility has experienced a power outage at approximately 1845 GMT.  Power was restored to the Central Facility at approximately 1945 GMT.  RTS @2000 GMT. 

7/17/02

MR

The instrument mentor wants the PSP instrument changed out.  He will send a spare.  The spare instrument arrived on site and needs to be installed.  Installed the *PSP.  RTS at 1900 GMT.  *Old - 960813  *New - 670956

7/15/02

MR

The instrument mentor wants the visible instrument connection checked for corrosion. Checked the instrument connection for corrosion.  I did not see any signs of corrosion but I cleaned the connection anyway.  Notified the mentor.  RTS at 1545 GMT. 

7/11/02

MR

The instrument mentor has sent us a replacement heater resistor/capacitor pair and wants it installed in the visible instrument control box.  Installed the components as per instructions sent along.  Notified the mentor that the procedure was performed.  RTS at 2045 GMT.

6/14/02

MR

The mentor notified me that the modem was locked up and would not answer for the visible instrument.  Power cycled the modem and the board.  The modem lights returned and seemed ok.  Notified the mentor.  RTS at 1630 GMT.

3/28/02

CAM

George Janson sent a replacement UVB-1 and requested that Site Ops exchange it with the sensor currently installed.  R/R *UVB-1 sensor at 1630 GMT.  RTS at 1640 GMT.  *Old – 930207  *New - 960813

7/16/01

CAM

Bill Durham contacted Site Ops and requested that the visible RSR data logger be power cycled to correct a 30mV offset that had developed.  Powered the data logger off at 1621 GMT.  Waited 10 minutes and powered back up at 1631 GMT.  Notified the mentor that the power cycle was complete and he logged on to the system to see if the problem was corrected.  At that time he found that the logger did not appear to reset and it was determined that the logger had to be cycled using the battery switch on the board and not the main power supply switch.   Powered off the battery switch at 1647 GMT and powered back on at 1657 GMT.  Mentor then reinitialized the logger remotely and advised that the problem had been corrected.

5/18/01

CAM

Bill Durham advised that the humidity sensor read 104% during recent rain and requested that an adjustment be made. Turned the dry adjustment screw 1/2 turn CCW as requested. RTS at 1615 GMT.

12/14/00

CAM

Received a new phone switch to be installed. R/R Command Communications (model ASAP 102) *phone switch with ComShare 450 phone switch. *Old - 1000102970600383 *New - 2000450970500066. Bill Durham checked and verified communications are now restored. RTS at 2100 GMT.

12/12/00

CAM

Bill Durham requested that the phone switch be bypassed so that the phone line connects directly to the UVRSR modem. This is only to confirm that the switch is faulty. Connected the phone line directly to the modem and Bill was then able to successfully log in and restart the UVRSR. Returned the phone line setup to its original configuration. Measured the phone line voltage as 46.9 VDC to make sure a new Com South switch would work. RTS at 1530 GMT. Bill will send a new switch to be installed.

12/11/00

MR

The mentor wanted the instrument modem power cycled. Power cycled the modem at 1510. The mentor said this did not solve the problem and wanted the instrument switch power cycled. Power cycled the switch at 1615 GMT.

8/7/00

MR

The replacement cable came in today and needs to be installed. Called Bill Durham and notified him. Installed the cable and called Bill back. He said the problem seems to still be there and he was going to call Yankee about it in the morning. Bill called me back at 2110 GMT and wanted me to power cycle the instrument just to make sure the data logger was not hung up. Power cycled the instrument at 2115 GMT. Will await further instruction from the mentor.

8/3/00

MR

Bill Durham sent a request that we do some testing for him over the phone. Called Bill at 1540 GMT. He had me remove the plug for the downwelling visible instrument and wait a few minutes to observe the data. He then had me unplug the sensor and allow it to sit for a few minutes. While doing the work, I noticed that the sensor cable was trying to pull from the connector. I notified him of this and he said he would send a new cable out and hoped this would solve the problem he was having with it. Plugged all the sensors and plugs back in and RTS at 1610 GMT.

6/5/2000

MR

Changed out the *UVB sensor head. Called and had data downloaded before change out. Called and had system started back up at 2022 GMT. Shading checked O.K. *Old - 00389 *New - 00299

3/29/2000

KT

Mentor request powering off the West side logger board for 15 minutes. At 2010 GMT the logger board was powered off. At 2026 GMT powered the logger board on. Status ok at this time. RTS at 2035 GMT.

12/22/99

MR

Changed the *data logger board. Replaced the old **MFR head. Called George Janson and had him reinitialize the board and adjust the shadow band. He reported that all the channels had returned. Old Board Sn# B444. Removed head Sn# 00388.

12/20/99

CAM

Mentor indicated that one channel of the UVRSR is not reading and shipped a new sensor to the site. Powered the UVRSR off at 1825 GMT. R/R *UVRSR sensor. Restored power at 1833 GMT. George Janson made adjustments to the shadow band stepper motor software to center the shadow on the diffuser. T/S found that the channel was still reporting zero when monitored remotely. Used a DVM to measure the signal being ouput by the sensor and found the sensor to be working properly. Measured the signal at several test points on the data logger and found the data logger is bad. Mentor will ship a replacement logger. Left the new sensor in place at this time. Suspect the original sensor is good and will confirm when replacement logger is received. *Old - 00389 *New - 00388

12/14/99

CAM

Technicians found during daily rounds on 12/13 that there was moisture inside the UVB-1 dome. Notified instrument mentor who advised he would send a replacement Pyranometer. Received a replacement UVB-1 on 12/14. R/R *UVB-1 Pyranometer. Removed old sensor at 2050 GMT. Installed new sensor, leveled, and RTS at 2055 GMT. RTS status O.K. Old sensor will be returned to mentor. *Old - 930810 (Sens 1.97 (W/m2)/V Cal 7/17/97) *New - 920903 (Sens 1.97 (W/m2)/V Cal 6/11/98)

10/21/99

MR

The mentor wants us to change out the logger board and the radiometer on the UVB1. Called George Janson at 1915 and he emptied the logger of data. R/R the *data logger and **UVB1 radiometer. Called George back and he started the instrument and looked at the readings and said the readings looked good. RTS at 2000. *Old - B98F *New - 93916 **Old - 687569 **New - 670957

7/12/99

MW

The instrument mentor called and said communications with the UVB had been lost over the previous weekend. Troubleshooting efforts revealed that the UVB modem had no LED's lit. Power cycled the modem at 1545 GMT. The LED's came back on. The mentor was able to communicate with the instrument after the power cycle and was able to download a data file from the logger. RTS at 1545 GMT.

4/8/99

CAM

Cleaned water spots from UVB1 dome at 1500 GMT.

4/7/99

CAM

Removed the failed Comshare450 *telephone switch and replaced with new ASAP102 switch at 1800 GMT. Verified proper output from the power transformer at 15.2 VAC. Reconnected the line in and MFRSR and UVRSR phone lines. George Janson was then able to collect data from each instrument going through the telephone switch. The surge suppresser also checked good. RTS at 1815 GMT. Will ship the failed switch back to George.

4/6/99

CAM

George Janson requested assistance to manually collect data from the UVRSR. The phone switch has not been working for several days now and UVRSR data is at risk of being lost. Connected the phone line directly to the UVRSR modem at 2115 GMT to allow mentor to collect data. Received a replacement switch today and will install tomorrow.

4/2/99

JEM

Weekly PM. Cleaned UVB1 dome and MFR sensors. Status O.K.

3/26/99

MK

Weekly PM. Status O.K.

3/19/99

MR

Weekly PM. Cleaned UVB1 dome and MFR sensors. Shading not confirmed due to ovc sky. RTS at 1430 GMT.

3/16/99

MW

The instrument mentor indicated a problem in communicating with both loggers. He asked that the Comshare switch be replaced. Replaced the switch and communications were re-established.

3/12/99

JEM

Weekly PM. Cleaned UVB1, MFR and UVMFR. Unable to check for proper shading due to ovc sky. Level O.K. RTS at 1535 GMT.

3/11/99

MR

PM. Cleaned the domes and the MFRSR heads.

3/10/99

MR

Status O.K.

3/9/99

MR

Status O.K.

3/8/99

MR

The UVMFR shadow band is hitting the side of the radiometer housing. Called Bill Durham and notified him. He suspected the motor control chip on the board and said there are 2 chips which one of them is a spare. He Had me power the unit down at 1600 and swap the main control chip with the spare one. Powered up the unit at 1615 and had Mr. Durham reinitialize the unit. This did not repair the problem. Power down the unit at 1619. Mr. Durham also wanted to check for corrosion in the motor plug socket. Removed the socket and found it was full of water. Cleaned and dried out the socket and plug with compressed air. Replaced the socket and called Mr. Durham and notified him of the problem. He Reinitialized the unit and it then properly worked. RTS the unit at 1700.

3/5/99

JM

Weekly PM. Cleaned dome at 1630 GMT.

3/3/99

MW

Cleaned dome at 1639 GMT.

2/26/99

KT

Weekly PM. Cleaned dome at 1440 GMT. Status O.K.

2/19/99

ZW

Weekly PM. Cleaned dome at 1439 GMT. Shadow band alignment could not be checked because of sky conditions. Level was ok. RTS 1445 GMT.

2/17/99

CAM

The USDA UV-B instrument array was successfully installed at the SGP CART.  The instruments were installed on the vacant Southern most mounting pole of the SIRS E13 instruments.  An analog phone line was installed an connected to the CF telephone distribution panel, and successful communications tests were performed with the USDA network control in Boulder, CO.